Suppose you are 94 cm from a plane mirror.

Part A

What area of the mirror is used to reflect the rays entering one eye from a point on the tip of your nose if your pupil diameter is 4.1 mm ? Express your answer to two significant figures and include the appropriate units.
Answer #1

Neglating the angle between nose and eye we have,

tan\theta = \frac{r}{d}

r= d tan \theta

r= 94\times tan\theta

Now calculating for the tangent will get,

As you are half way between your image and you, we have to consider tanget as,

tan\theta = \frac{0.41}{4\times 94}

tan\theta = 1.09\times 10^{-3}

\therefore r =102.46 \times 10^{-3}cm

Area is given by

A=\pi r^2

A=3.14\times {(102.46\times 10^{-5})}^2

A=3.30\mu m